An auto's velocity increases uniformly from 6.0 m/s to 20 m/s while covering 70 m in a straight line. Find the acceleration and the time taken.


1
Expert's answer
2021-03-29T08:58:50-0400

(a)

"v^2=v_0^2+2as,""a=\dfrac{v^2-v_0^2}{2s},""a=\dfrac{(20\ \dfrac{m}{s})^2-(6\ \dfrac{m}{s})^2}{2\cdot70\ m}=2.6\ \dfrac{m}{s^2}."

(b)

"v=v_0+at,""t=\dfrac{v-v_0}{a},""t=\dfrac{20\ \dfrac{m}{s}-6\ \dfrac{m}{s}}{2.6\ \dfrac{m}{s^2}}=5.4\ s."